find the value of y so that the slope of the line passing through
(-3,-12) and (1,y) is 2

Answer by tarungehlot(9)

for any two point (x1,y1) and (x2,y2)
slope= y2 -y1/x2-x1
2= y+12/1+3
2= y+12/4
y+12=8
y=-4
hence
points are
(-3 ,12 ) and ( 1,-4)